OpenStudy (anonymous):

PLEASE HELP! Sketch theta = 5pi/4 in standard position. A. Determine two coterminal angles (pos & neg) B. Convert the angle to degree measure.

OpenStudy (anonymous):

It may help to convert 5pi/4 into degrees

OpenStudy (anonymous):

(5*180)/4 = 225 degrees, so it is in quad 3

OpenStudy (anonymous):

so 225 degrees would be the conversion of the angle to its degree measurement?

OpenStudy (anonymous):

would that also be one of the coterminal angles?

OpenStudy (anonymous):



OpenStudy (anonymous):

Terminal angles is the angle + or - 360 degrees. sin(225 degrees + 360 degrees) = sin(585 degrees) = sin(225 degrees)

OpenStudy (anonymous):

In the sketch, you would label it as 5pi/4 rads

OpenStudy (anonymous):

so the two coterminal angles would be -135 and 585 degrees, the conversion of the angle to degree measure would be 225 degrees and the sketch would be a horizontal line at positive 4, correct?

OpenStudy (anonymous):

two terminal angles of 225 degrees are: (225-360) degrees, and (225+360) degrees

OpenStudy (anonymous):

what do you mean by horizontal line at positive four? The unit circles goes like this: http://coolmath.com/precalculus-review-calculus-intro/precalculus-trigonometry/images/28-trigonometry-01.gif Do you see where 5pi/4 is?

OpenStudy (anonymous):

Other than that, everything else is correct

OpenStudy (anonymous):

thank you!!! :)